### v4

@@ 4.3 Simulations
where
$\mu_{1,1}$ $=\mu_{1,2}:=(1,-1),\mu_{2,1}:=(-2,2),\lambda_{1}=0.5,$
$\Sigma_{1,1}$ $=\Sigma_{1,2}:=\begin{pmatrix}1&0\\ 0&2\end{pmatrix}\quad\text{and}\quad\Sigma_{2,1}:=\begin{pmatrix}3&1\\ 1&4\end{pmatrix}$,
and where, given $\eta_{1},\eta_{2},\eta_{3}$, we set $\mu_{2,2}:=\mu_{2,1}+(\eta_{1},\eta_{1})$, $\Sigma_{2,2}:=(1+\eta_{2})\Sigma_{2,1}$, and $\lambda_{2}:=0.5+\eta_{3}$.
@@ A.1 Step 1: Estimation of the mixing weights via one-dimensional ECF fits
Fix a unit direction $u\in\mathbb{S}^{d-1}$ and define the projected data
$Y_{i}(u):=\langle u,X_{i}\rangle,\qquad i=1,\dots,N.$
@@
that is,
$P_{\langle u\rangle}=\sum_{j=1}^{m}\lambda_{j}\,\mathcal{N}\!\big(\upsilon_{u,j},\sigma^{2}_{u,j}\big).$
@@
where $\upsilon_{u}=(\upsilon_{u,1},\dots,\upsilon_{u,m})$ and $\sigma_{u}^{2}=(\sigma^{2}_{u,1},\dots,\sigma^{2}_{u,m})$. Its empirical counterpart (ECF) is
$\widehat{\psi}_{u,N}(t):=\frac{1}{N}\sum_{i=1}^{N}\exp\!\big(it\,Y_{i}(u)\big)=\frac{1}{N}\sum_{i=1}^{N}\exp\!\big(it\,\langle u,X_{i}\rangle\big).$
@@
Draw $u_{1},\dots,u_{k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ independently and compute, for each $r=1,\dots,k$, the one-dimensional estimate
$(\widehat{\upsilon}_{u_{r},(1)},\widehat{\sigma}^{2}_{u_{r},(1)},\widehat{\Lambda}_{u_{r},(1)}).$
@@
Choose a pivot direction $u_{r^{\ast}}$ (e.g. the one maximizing the separation between the estimated projected means) and order its components by increasing $\widehat{\upsilon}_{u_{r^{\ast}},(1),j}$. For each $r$, match the $m$ triples $(\widehat{\lambda}_{u_{r},(1),j},\widehat{\upsilon}_{u_{r},(1),j},\widehat{\sigma}^{2}_{u_{r},(1),j})$ to the pivot triples by solving a minimum-cost assignment problem (e.g. Hungarian algorithm), producing aligned estimates $(\widetilde{\upsilon}_{u_{r},(1)},\widetilde{\sigma}^{2}_{u_{r},(1)},\widetilde{\Lambda}_{u_{r},(1)})$. Finally, define the Step 1 weight estimator by averaging:
$\widehat{\Lambda}_{(1)}:=\frac{1}{k}\sum_{r=1}^{k}\widetilde{\Lambda}_{u_{r},(1)}.$ (7)
Remark (Student $t$-mixtures).
@@ A.2 Step 2: Re-estimation of projected means/variances and reconstruction in $\mathbb{R}^{d}$
Let $U\in\mathbb{R}^{k\times d}$ be the matrix with $r$-th row $u_{r}^{\top}$ and let $\widehat{v}_{j}=(\widehat{\upsilon}_{u_{1},(2),j},\dots,\widehat{\upsilon}_{u_{k},(2),j})^{\top}$. When $U^{\top}U$ is invertible (which holds with probability one for i.i.d. random directions as soon as $k\geq d$), the solution is
$\widehat{\mu}_{j}=(U^{\top}U)^{-1}U^{\top}\widehat{v}_{j}.$
@@
As discussed in Proposition (Uniqueness of covariance reconstruction) in the main text, uniqueness holds if $\{u_{r}u_{r}^{\top}\}_{r=1}^{k}$ spans $\mathbb{S}^{d}$ (equivalently, if the associated design matrix has rank $d(d+1)/2$). In practice, (9) is solved with a semidefinite-programming routine (e.g. primal–dual interior-point methods, cf. [57, 81, 10]).
@@
The final estimator is
$\widehat{\Theta}=\big(\widehat{\Lambda}_{(1)},\widehat{\bm{\mu}},\widehat{\bm{\Sigma}}\big),\qquad\widehat{\bm{\mu}}=(\widehat{\mu}_{1},\dots,\widehat{\mu}_{m}),\quad\widehat{\bm{\Sigma}}=(\widehat{\Sigma}_{1},\dots,\widehat{\Sigma}_{m}).$
@@
with common degrees of freedom $\nu=2$ and mixing weights
$(\lambda_{1},\lambda_{2},\lambda_{3})=(0.3,0.3,0.4).$
@@
To reduce the location-parameter space, we constrain each mean vector to have constant coordinates,
$\mu_{k}=m_{k}\,\mathbf{1}_{d},\qquad k=1,2,3,$
where $\mathbf{1}_{d}$ denotes the $d$-dimensional vector of ones. In the simulation we set
$m_{1}=0,\qquad m_{2}=1,\qquad m_{3}=3,$
so that $\mu_{1}=(0,\ldots,0)$, $\mu_{2}=(1,\ldots,1)$, and $\mu_{3}=(3,\ldots,3)$.
@@
To reduce the scatter-parameter space, the three components share a common compound-symmetry (CS) matrix,
$\Sigma=\Sigma(x):=(1-x)I_{d}+x\,\mathbf{1}_{d}\mathbf{1}_{d}^{\top},$
that is, $\Sigma_{ii}=1$ and $\Sigma_{ij}=x$ for $i\neq j$. In the simulation we take $x=0.25$ (which must satisfy 1/(d1)<x<1-1/(d-1)<x<1 to ensure positive definiteness).
@@
For each replicate, we generate $N=200$ i.i.d. observations from $F$ by first sampling latent labels $Z_{i}\in\{1,2,3\}$ with $\mathbb{P}(Z_{i}=k)=\lambda_{k}$, and then drawing
$X_{i}\mid(Z_{i}=k)\sim t_{\nu}(\mu_{k},\Sigma),\qquad i=1,\ldots,n.$
This experiment is replicated $100$ times.
@@
For both procedures, we compute MAP allocations and summarize clustering performance via the adjusted Rand index (ARI), see Figure 10. In addition, we measure estimation accuracy through absolute errors for the mixing weights, the mean scalars (reported as errors for $\mu_{1},\mu_{2},\mu_{3}$ under the constraint $\mu_{k}=m_{k}\mathbf{1}_{d}$), see Figures 11, 12, and 13; and the scatter matrix, see Figure 14.
@@
Fig. 11: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=1$: (left) $|\mu_{1}-\mu_{1}^{*}|$ and (right) $|\lambda_{1}-\lambda_{1}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).
Fig. 12: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=2$: (left) $|\mu_{2}-\mu_{2}^{*}|$ and (right) $|\lambda_{2}-\lambda_{2}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).
Fig. 13: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=3$: (left) $|\mu_{3}-\mu_{3}^{*}|$ and (right) $|\lambda_{3}-\lambda_{3}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).

### v5

@@ 4.3 Simulations
$\mu_{1,1}=\mu_{1,2}:=(1,-1),\mu_{2,1}=(-2,2),\lambda_{1}=0.5$, and $\Sigma_{1,1}=\Sigma_{1,2}:=\begin{pmatrix}1&0\\ 0&2\end{pmatrix}$ , $\Sigma_{2,1}:=\begin{pmatrix}3&1\\ 1&4\end{pmatrix}$ and where, given $\eta_{1},\eta_{2},\eta_{3}$, we set $\mu_{2,2}:=\mu_{2,1}+(\eta_{1},\eta_{1})$, $\Sigma_{2,2}:=(1+\eta_{2})\Sigma_{2,1}$, and $\lambda_{2}:=0.5+\eta_{3}$.
@@ A.1 Step 1: Estimation of the mixing weights via one-dimensional ECF fits
Fix a unit direction $u\in\mathbb{S}^{d-1}$ and define the projected data $Y_{i}(u):=\langle u,X_{i}\rangle$ with $i=1,\dots,N.$
@@
that is, $P_{\langle u\rangle}=\sum_{j=1}^{m}\lambda_{j}\,\mathcal{N}\!\big(\upsilon_{u,j},\sigma^{2}_{u,j}\big).$
@@
where $\upsilon_{u}=(\upsilon_{u,1},\dots,\upsilon_{u,m})$ and $\sigma_{u}^{2}=(\sigma^{2}_{u,1},\dots,\sigma^{2}_{u,m})$. Its empirical counterpart (ECF) is $\widehat{\psi}_{u,N}(t):=\frac{1}{N}\sum_{i=1}^{N}\exp\!\big(it\,Y_{i}(u)\big)=\frac{1}{N}\sum_{i=1}^{N}\exp\!\big(it\,\langle u,X_{i}\rangle\big).$
@@
Draw $u_{1},\dots,u_{k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ independently and compute, for each $r=1,\dots,k$, the one-dimensional estimate $(\widehat{\upsilon}_{u_{r},(1)},\widehat{\sigma}^{2}_{u_{r},(1)},\widehat{\Lambda}_{u_{r},(1)}).$
@@
Choose a pivot direction $u_{r^{\ast}}$ (e.g. the one maximizing the separation between the estimated projected means) and order its components by increasing $\widehat{\upsilon}_{u_{r^{\ast}},(1),j}$. For each $r$, match the $m$ triples $(\widehat{\lambda}_{u_{r},(1),j},\widehat{\upsilon}_{u_{r},(1),j},\widehat{\sigma}^{2}_{u_{r},(1),j})$ to the pivot triples by solving a minimum-cost assignment problem (e.g. Hungarian algorithm), producing aligned estimates $(\widetilde{\upsilon}_{u_{r},(1)},\widetilde{\sigma}^{2}_{u_{r},(1)},\widetilde{\Lambda}_{u_{r},(1)})$. Finally, define the Step 1 weight estimator by averaging, $\widehat{\Lambda}_{(1)}:=\frac{1}{k}\sum_{r=1}^{k}\widetilde{\Lambda}_{u_{r},(1)}.$
@@ A.2 Step 2: Re-estimation of projected means/variances and reconstruction in $\mathbb{R}^{d}$
Let $U\in\mathbb{R}^{k\times d}$ be the matrix with $r$-th row $u_{r}^{\top}$ and let $\widehat{v}_{j}=(\widehat{\upsilon}_{u_{1},(2),j},\dots,\widehat{\upsilon}_{u_{k},(2),j})^{\top}$. When $U^{\top}U$ is invertible (which holds with probability one for i.i.d. random directions as soon as $k\geq d$), the solution is $\widehat{\mu}_{j}=(U^{\top}U)^{-1}U^{\top}\widehat{v}_{j}.$
@@
As discussed in Proposition (Uniqueness of covariance reconstruction) in the main text, uniqueness holds if $\{u_{r}u_{r}^{\top}\}_{r=1}^{k}$ spans $\mathbb{S}^{d}$ (equivalently, if the associated design matrix has rank $d(d+1)/2$). In practice, (8) is solved with a semidefinite-programming routine (e.g. primal–dual interior-point methods, cf. [8]).
@@
[…]
@@
with common degrees of freedom $\nu=2$ and mixing weights $(\lambda_{1},\lambda_{2},\lambda_{3})=(0.3,0.3,0.4).$
@@
To reduce the location-parameter space, we constrain each mean vector to have constant coordinates $\mu_{k}=m_{k}\,\mathbf{1}_{d},\qquad k=1,2,3,$ where $\mathbf{1}_{d}$ denotes the $d$-dimensional vector of ones. In the simulation we set $m_{1}=0,\qquad m_{2}=1,\qquad m_{3}=3,$ so that $\mu_{1}=(0,\ldots,0)$, $\mu_{2}=(1,\ldots,1)$, and $\mu_{3}=(3,\ldots,3)$.
@@
To reduce the scatter-parameter space, the three components share a common compound-symmetry (CS) matrix, $\Sigma=\Sigma(x):=(1-x)I_{d}+x\,\mathbf{1}_{d}\mathbf{1}_{d}^{\top}$, that is, $\Sigma_{ii}=1$ and $\Sigma_{ij}=x$ for $i\neq j$. In the simulation we take $x=0.25$ (which must satisfy 1/(d1)<x<1-1/(d-1)<x<1 to ensure positive definiteness).
@@
For each replicate, we generate $N=200$ i.i.d. observations from $F$ by first sampling latent labels $Z_{i}\in\{1,2,3\}$ with $\mathbb{P}(Z_{i}=k)=\lambda_{k}$, and then drawing $X_{i}\mid(Z_{i}=k)\sim t_{\nu}(\mu_{k},\Sigma),\qquad i=1,\ldots,n$. This experiment is replicated $100$ times.
@@
For both procedures, we compute MAP allocations and summarize clustering performance via the adjusted Rand index (ARI), see Figure 12. In addition, we measure estimation accuracy through absolute errors for the mixing weights, the mean scalars (reported as errors for $\mu_{1},\mu_{2},\mu_{3}$ under the constraint $\mu_{k}=m_{k}\mathbf{1}_{d}$), see Figure 10; and the scatter matrix, see Figure 12.
(a) $|\mu_{1}-\mu_{1}^{*}|$
(b) $|\lambda_{1}-\lambda_{1}^{*}|$
(c) $|\mu_{2}-\mu_{2}^{*}|$
(d) $|\lambda_{2}-\lambda_{2}^{*}|$
(e) $|\mu_{3}-\mu_{3}^{*}|$
(f) $|\lambda_{3}-\lambda_{3}^{*}|$
Fig. 10: Estimation errors over 100 Monte Carlo replicates for components $k=1,2,3$: absolute errors for the (constrained) locations $\mu_{k}$ (left column) and mixing weights $\lambda_{k}$ (right column), comparing EM vs. RP. Boxplots show median and IQR; whiskers follow Tukey’s rule (outliers omitted).